# Supercritical fluids as a distinct state of matter characterized by sub-short-range structural order

### v1

@@ Methods
where $\epsilon/k_{\rm B}=119.5$ K, $\sigma=$3.405\text{\,}\mathrm{\text{\AA }}$\$, and $k_{\rm B}=$1.38\text{\times}{10}^{-23}\text{\,}\mathrm{J}\text{\,}{\mathrm{K}}^{-1}$\$ is the Boltzmann constant. The initial system is a face-centered cubic lattice of 4000 atoms with a lattice constant of $5.26\text{\,}\mathrm{\text{\AA }}$. Thirty-five different pressures were simulated from $5\text{\,}\mathrm{bar}$ to $1300\text{\,}\mathrm{bar}$. For each target pressure, the system is initially relaxed at $195\text{\,}\mathrm{K}$ for $5\text{\,}\mathrm{ns}$ in an isothermal-isobaric (NPT) ensemble, in order to reach thermal equilibrium. Then the simulation is performed in a canonical (NVT) ensemble for $20\text{\,}\mathrm{ns}$ – all measurements are performed in this time period. Temperature and pressure are regulated via the Nosé-Hoover methods Nosé (1984); Hoover (1985), with Newton’s equations of motion solved numerically using a time step of $1\text{\,}\mathrm{fs}$. All samples are simulated with periodic boundary conditions. The pressure unit is bar, length unit Angstroms, time unit fs, and energy unit kcal/mol.
@@ Acknowledgements.
SJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). XL acknowledges the Postdoctoral Fellowship Program of CPSF (Grant Number GZC20252776). XF acknowledges the support of the National Natural Science Foundation of China (No.52201016). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. YJ acknowledges funding from NSFC (Grant 12047503) and Wenzhou Institute (Grant WIUCASICTP2022). SJ and XF thank the Key State Laboratory for Clean and Efficient Turbomachinery Power Equipment for providing HPC resources. The authors acknowledge the use of the High Performance Cluster at Institute of Theoretical Physics, Chinese Academy of Sciences.
Supplementary Material
@@
In the supplementary material we provide extended data, details of the analyses presented in the main text and additional technical information about the theoretical frameworks used.
@@
Figure S1: Extended data. (a) Radial distribution functions. Data for argon are obtained from MD simulations at P=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800barP=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800\ $\text${bar} (from red to blue) and T=195KT=195\ $\text${K}. (b) Velocity auto-correlation functions. Data for argon are obtained from MD simulations at P=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300barP=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300\ $\text${bar} (from right to left) and T=195KT=195\ $\text${K}.
Figure S2: Estimate of the SSR length $\xi$. Radial distribution functions around the (a) third ($n=3$) and (b) fourth ($n=4$) peaks. (inset) The height of the third ($h_{3}$) and fourth ($h_{4}$) peaks as a function of $P$.
Figure S3: Evaluation of the length scale $r^{*}$. Here $r^{*}$ is obtained by an exponential fit of the large-$r$ data of $\tilde{h}(r)=r[g(r)-1]$. (a) For $P<P^{+}=285$ bar (below the $L^{+}$ line), the fitting is performed directly to $\tilde{h}(r)$. (b) For $P>P^{\rm+}$ (above the $L^{+}$ line), the fitting is performed to the envelope of peaks. Note that when $P<P^{+}$, there are only a few peaks at small $r$; the tail of $\tilde{h}(r)$ does not show peaks.
Figure S4: SSR structural order and characteristic lengths along the isothermal line T=220KT=220\ $\text${K}. Red line and blue line represent the SSR length $\xi$ and asymptotic decay length $r^{*}\equiv\max(r_{0},r_{1})$ as functions of $P$ respectively.
Figure S5: Effective length unit from the viscosity data. We assume that the kinetic expression of viscosity can be generalized to higher pressures, with a variable effective particle diameter $\sigma_{\rm eff}(P)$, $\eta=\frac{5}{16\sigma_{\rm eff}^{2}(P)}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2}$, where $\sigma_{\rm eff}$ is obtained by fitting the NIST data of $\eta$. The growth of $\sigma_{0}/\sigma_{\rm eff}$ begins around the $L^{-}$ line.
Figure S6: Reduced diffusion coefficient as a function of the two-particle structural entropy $S_{2}$. The solid line represents an exponential fit of the data above the $L^{+}$ line, $D=1.18e^{S_{2}/k_{\rm B}}$.
Figure S7: Einstein frequency $\Omega$ (Thz) as a function of pressure $P$. The power-law fits (lines) are guides to the eye.
Theories of auto-correlation functions and transport coefficients
Velocity auto-correlation function (VACF) and diffusion coefficient
@@
Lucks-Moser theory
@@
Stress auto-correlation function and Viscosity
@@
Lucks-Moser theory

### v2

@@ Methods
where $\epsilon/k_{\rm B}=119.5$ K, $\sigma=$3.405\text{\,}\mathrm{\text{\AA }}$\$, and $k_{\rm B}=$1.38\text{\times}{10}^{-23}\text{\,}\mathrm{J}\text{\,}{\mathrm{K}}^{-1}$\$ is the Boltzmann constant. To address potential finite-size effects, we have evaluated systems of 500, 4,000, and 32,000 atoms. Our analysis confirm that 4,000 atoms provides a balance between statistical accuracy and computational efficiency for calculating transport properties; consequently, this system size is adopted for the production runs. The initial system is a face-centered cubic lattice of 4000 atoms with a lattice constant of $5.26\text{\,}\mathrm{\text{\AA }}$. Thirty-five different pressures are simulated from $5\text{\,}\mathrm{bar}$ to $1300\text{\,}\mathrm{bar}$. For each target pressure, the system is initially relaxed at $195\text{\,}\mathrm{K}$ for $5\text{\,}\mathrm{ns}$ in an isothermal-isobaric (NPT) ensemble to reach thermal equilibrium. Following the NPT stage, the system undergoes an intermediate 500 ps relaxation in the microcanonical (NVE) ensemble at each specific state point to ensure total energy stability and facilitate unconstrained dynamical relaxation. Finally, a canonical (NVT) relaxation is performed to stabilize the system prior to the $20\text{\,}\mathrm{ns}$ production phase – all measurements are performed in this time period. The temperature and pressure are regulated via the Nosé-Hoover methods Nosé (1984); Hoover (1985), with Newton’s equations of motion solved numerically using a time step of $1\text{\,}\mathrm{fs}$. All samples are simulated with periodic boundary conditions. The pressure unit is bar, length unit Angstroms, time unit fs, and energy unit kcal/mol.
@@ Acknowledgements.
We thank S.Khrapak for bringing the recent Ref. Khrapak et al. (2025) to our attention. SJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). XL acknowledges the Postdoctoral Fellowship Program of CPSF (Grant Number GZC20252776). XF acknowledges the support of the National Natural Science Foundation of China (No.52201016). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. YJ acknowledges funding from the National Key R&D Program of China (Grant No. 2025YFF0512000), Project supported by the Space Application System of China Manned Space Program, NSFC (Grant NO. 12447101) and Wenzhou Institute (Grant WIUCASICTP2022). SJ and XF thank the Key State Laboratory for Clean and Efficient Turbomachinery Power Equipment for providing High Performance Cluster resources. The authors acknowledge the use of the High Performance Cluster at Institute of Theoretical Physics, Chinese Academy of Sciences, and the computer clusters at the Hefei advanced computing center.
Supplementary Information
@@
In the supplementary information we provide extended data, details of the analyses presented in the main text and additional technical information about the theoretical frameworks used.
S1 Supplementary analysis of radial distribution functions (RDFs).
In Fig. S1, we plot a complete set of RDFs measured by molecular dynamics (MD) simulations at $T=195$ K and different pressures. The radial distribution functions are calculated by averaging over a large number of frames to reduce statistical noise. We extract 500 ps of trajectories with 1 fs intervals, resulting in 500,000 frames for averaging. These extensive averages ensure that the RDFs are well-converged.
In Fig. S2, we show RDFs enlarged around the second peak, at another temperature, $T=222$ K. The second peak emerges around $P_{2}=55$ bar, consistent with the pressure $P^{-}=47$ bar at the $L^{-}$ line. This test provides an additional verification of the structural signature at the $L^{-}$ line.
Next, we discuss the procedure employed to determine the peaks in $g(r)$. To quantitatively track the evolution of structural features in the radial distribution function $g(r)$, its second peak is fitted using a Gaussian function. Specifically, we employ a Gaussian function of the following form to fit the second peak,
$f(r)=c+h_{2}\exp\!\left[-\frac{(r-r_{0})^{2}}{2\sigma^{2}}\right],$ (S1)
with fitting parameters $c,h_{2},r_{0}$ and $\sigma$. Here $h_{2}$ and $r_{0}$ provide the height of position respectively. When the coefficient $R^{2}$ of determination is below $0.8$, the Gaussian model no longer provides an adequate description of the feature, indicating that the second peak cannot be reliably identified.
@@
Figure S1: Radial distribution functions at T=195KT=195\ $\text${K}. Radial distribution functions $g(r)$ for argon are obtained from MD simulations at P=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800barP=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800\ $\text${bar} (from red to blue) and T=195KT=195\ $\text${K}.
Figure S2: Enlarged radial distribution functions at T=220KT=220\ $\text${K}. Radial distribution function $g(r)$ at different pressures, zoomed in around the second peak, for T=220KT=220\ $\text${K}.
S2 Theories of auto-correlation functions and transport coefficients
S2.1 Velocity auto-correlation function (VACF) and diffusion coefficient
@@
S2.1.2 Lucks-Moser theory
@@
S2.2 Stress auto-correlation function and Viscosity
@@
S2.2.2 Lucks-Moser theory
@@
S3 Supplementary analysis of velocity auto-correlation functions (VACFs)
In Fig.S3, we plot a complete set of VACFs measured by molecular dynamics (MD) simulations at $T=195$ K and different pressures.
Next, we detail the fitting procedure used in Fig. 4a. The VACF is computed for all simulated pressures and fitted over a time window chosen to ensure convergence while minimizing numerical noise from the long-time tail. Because the convergence rate of the VACF strongly depends on pressure—being slower at low pressures and faster at high pressures—the fitting window for the stretched-exponential form is adjusted accordingly. For instance, at $5$ bar the VACF is evaluated over $100$ ps and fitted in ten consecutive $5$ ps segments starting from $5$ ps. As pressure increases and correlations decay more rapidly, progressively shorter fitting windows are used: $80$ ps at $10$ bar, $60$ ps at $15$ bar, $50$ ps at 20202525 bar, and 4466 ps for pressures above $200$ bar. For the highest-pressure states ($\geq 400$ bar), the VACF converges within $4$ ps, and the fitting is performed in ten consecutive $0.1$ ps segments starting from $0.8$ ps.
Figure S3: Velocity autocorrelation functions. Velocity auto-correlation functions $Z(t)$ for argon are obtained from MD simulations at P=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300barP=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300\ $\text${bar} (from right to left) and T=195KT=195\ $\text${K}.
S4 Calculation of $S(q)$ and Ornstein-Zernike Correlation Length from MD data.
The static structure factor $S(q)$ is computed directly from molecular dynamics trajectories. For each configuration snapshot containing $N=4000$ argon atoms, $S(\mathbf{q})$ is calculated as:
$S(\mathbf{q})=\frac{1}{N}\left|\sum_{j=1}^{N}e^{-i\mathbf{q}\cdot\mathbf{r}_{j}}\right|^{2},$ (S35)
where $\mathbf{r}_{j}$ are atomic positions and $\mathbf{q}$ are wave vectors compatible with the periodic boundary conditions. The wave vectors are grouped into spherical shells of width $\Delta q=0.05$ Å-1, and $S(\mathbf{q})$ values within each shell are averaged to obtain $S(q)$. For each pressure, $S(q)$ is evaluated over 500 ps and averaged over 500,000 configurations.
The Ornstein-Zernike (OZ) correlation length is obtained by fitting the low-$q$ region of $S(q)$ to the OZ form:
$S(q)=\frac{S_{0}}{1+\xi_{\mathrm{OZ}}^{2}q^{2}}.$ (S36)
Inverting this relation, one gets
$\frac{1}{S(q)}=\frac{1}{S_{0}}+\frac{\xi_{\mathrm{OZ}}^{2}}{S_{0}}q^{2},$ (S37)
from which $\xi_{\mathrm{OZ}}$ is determined via linear regression. Uncertainties are estimated from the fitting errors.
Figure S4: Ornstein-Zernike correlation length. $\xi_{\mathrm{OZ}}$ as a function of pressure at $T=195$ K, showing a maximum near the Widom line between the $L^{+}$ and $L^{-}$ lines. Error bars represent fitting uncertainties.
The resulting $\xi_{\mathrm{OZ}}$ as a function of pressure is shown in Fig. S4. $\xi_{\mathrm{OZ}}$ reaches its maximum near the Widom line, which is located roughly in between the $L^{\pm}$ lines.
S5 Identification of the Widom delta
Following the approach of Yoon et al. Jun et al. (2018), we employ a probabilistic classification algorithm to characterize the gas-like and liquid-like domains in the simulated argon systems.
For each configuration from MD data, we perform the Voronoi tessellation. The Voronoi cell of atom $i$ is defined as the set of points closer to atom $i$ than to any other atom. The local density $\rho_{i}^{\text{local}}$ is then calculated as the inverse of the Voronoi cell volume:
$\rho_{i}^{\text{local}}=\frac{1}{V_{i}^{\text{Voronoi}}},$ (S38)
where $V_{i}^{\text{Voronoi}}$ is the volume of the Voronoi polyhedron associated with atom $i$.
To account for spatial correlations and reduce sensitivity to local fluctuations, we compute the mean local density $\bar{\rho}_{i}$ for each atom $i$ by averaging over its Voronoi neighbors:
$\bar{\rho}_{i}=\frac{1}{N_{\text{neigh}}+1}\left(\rho_{i}^{\text{local}}+\sum_{j\in\mathcal{N}(i)}\rho_{j}^{\text{local}}\right),$ (S39)
where $\mathcal{N}(i)$ denotes the set of Voronoi neighbors of atom $i$ (atoms sharing a Voronoi face), and $N_{\text{neigh}}=|\mathcal{N}(i)|$ is the number of such neighbors.
The classification of atoms into gas-like and liquid-like states follows a probabilistic algorithm inspired by the work of Yoon et al. Jun et al. (2018):
  1. 1.
    Random processing order: Atoms are processed in random order. When an atom is selected, the mean local density of that atom and all its Voronoi neighbors is computed. This neighborhood averaging reduces sensitivity to local fluctuations.
  2. 2.
    Threshold comparison: For the current unlabeled atom $i$, its mean local density $\bar{\rho}_{i}$ is compared to a threshold value $\rho_{\text{th}}$. We use the critical density of argon $\rho_{c}$ as the threshold:
    $\text{label}=\begin{cases}1&(\text{liquid-like})\quad\text{if }\bar{\rho}_{i}\geq\rho_{c},\\ 0&(\text{gas-like})\quad\text{if }\bar{\rho}_{i}<\rho_{c}.\end{cases}$ (S40)
  3. 3.
    Simultaneous labeling: The chosen atom $i$ and all its unlabeled Voronoi neighbors are assigned the same label, reflecting the correlated nature of local environments.
  4. 4.
    Multiple trials: Steps 1–3 are repeated $N_{\text{trials}}=300$ times with different random processing orders to obtain statistically robust results.
  5. 5.
    Probability calculation: The liquid-like probability $P_{i}^{\text{liq}}$ for atom $i$ is calculated as the fraction of trials in which it is classified as liquid-like:
    $P_{i}^{\text{liq}}=\frac{1}{N_{\text{trials}}}\sum_{t=1}^{N_{\text{trials}}}L_{i}^{(t)},$ (S41)
    where $L_{i}^{(t)}=1$ if atom $i$ is labeled liquid-like in trial $t$, and $L_{i}^{(t)}=0$ otherwise.
For each pressure, we calculate the fractions of gas-like and liquid-like atoms as ensemble averages. The gas-like fraction $\Pi_{\text{gas}}$ is defined as:
$\Pi_{\text{gas}}=\frac{1}{N_{\text{atoms}}}\sum_{i=1}^{N_{\text{atoms}}}\mathcal{I}\left(P_{i}^{\text{liq}}\leq 0.5\right),$ (S42)
where $\mathcal{I}(\cdot)$ is the indicator function. Similarly, the liquid-like fraction $\Pi_{\text{liq}}$ is:
$\Pi_{\text{liq}}=1-\Pi_{\text{gas}}.$ (S43)
The fractions are averaged over 100 random configurations along the trajectories. Uncertainties in these fractions are estimated from the standard deviation across multiple simulation frames.
The resulting fractions of gas-like and liquid-like atoms as functions of pressure for simulated argon at two temperatures is shown in Fig.S5. The gas-like fraction $\Pi_{\text{gas}}$ approaches unity near the $L^{-}$ line, while the liquid-like fraction $\Pi_{\text{liquid}}$ approach unity much earlier than the $L^{+}$ line.
The Widom delta is defined from the inverse sigmoidal fit to $\Pi_{\text{gas}}$ by finding the densities (pressures) where the tangent line at $\Pi_{\text{gas}}=0.5$ intersects $\Pi_{\text{gas}}=1.0$ and $\Pi_{\text{gas}}=0.0$, respectively. We find that the Widom delta is within the intermediate transition regime bounded by $L^{\pm}$ lines.
Figure S5: Comparison between Widom delta and $L^{\pm}$ lines.Fractions of gas-like ($\Pi_{\text{gas}}$) and liquid-like ($\Pi_{\text{liquid}}$) atoms as functions of pressure for simulated argon at two temperatures: $\mathbf{(a)}$ $T=195\,\mathrm{K}$ and $\mathbf{(b)}$ $T=220\,\mathrm{K}$. The red dashed line indicates the Widom line.